sqrt(2x-5) - sqrt(x-3) = 1
Answer by mukhopadhyay(490)

sqrt(2x-5) - sqrt(x-3) = 1
=> sqrt(2x-5) = 1 + sqrt(x-3)
=> ]sqrt(2x-5)]^2 = [1 + sqrt(x-3)]^2
=> 2x-5 = 1+x-3+2sqrt(x-3)
=> x-3 = 2sqrt(x-3)
=> [x-3]^2 = [2sqrt(x-3)]^2
=> [x-3]^2 = 4(x-3)
=> (x-3)(x-3-4) = 0
=> (x-3)(x-7) = 0
=> x=3 or x=7